### v1

# Suppressing spurious transitions using spectrally balanced pulse

Abstract
Achieving precise control over quantum systems presents a significant challenge, especially in many-body setups, where residual couplings and unintended transitions undermine the accuracy of quantum operations. In superconducting qubits, parasitic interactions—both between distant qubits and with spurious two-level systems—can severely limit the performance of quantum gates. In this work, we introduce a pulse-shaping technique that uses spectrally balanced microwave pulses to suppress undesired transitions. Experimental results demonstrate an order-of-magnitude reduction in spurious excitations between weakly detuned qubits, as well as a substantial decrease in single-qubit gate errors caused by a strongly coupled two-level defect over a broad frequency range. Our method provides a simple yet powerful solution to mitigate adverse effects from parasitic couplings, enhancing the fidelity of quantum operations and expanding feasible frequency allocations for large-scale quantum devices.
Introduction — Quantum information processing technologies have made significant progress, achieving coherent control over systems with about 100 qubits [1, 2, 3, 4, 5, 6]. In superconducting quantum processors, qubit-qubit interactions are typically mediated by engineered coupling elements, such as coplanar capacitors. However, due to the long-range nature of electromagnetic interactions and design constraints, parasitic or residual couplings can exist between qubits that are intended to be uncoupled. These parasitic interactions, known as quantum crosstalk [7], pose a challenge to the execution of independent operations, as they can interfere with transitions close in frequency. This phenomenon degrades the performance of quantum operations, limiting the scalability of quantum computing systems.
Compounding this issue is the presence of uncontrolled microscopic degrees of freedom, often referred to as two-level defects or two-level systems (TLS), which are ubiquitous in many quantum platforms [8]. When frequencies of TLSs come close to those of the qubits, they can significantly interfere with qubit operations. Excitations into a long-lived TLS can be particularly harmful as they can accumulate over time and deteriorate subsequent operations, causing correlated or non-Markovian errors [9]. Without reset capability, this may be a potential threat to quantum error correction [10]. Moving the qubit frequency away from these TLSs can reduce their impact. However, this approach is not directly available for fixed-frequency qubits. Even for tunable qubits, the available frequency options are often severely constrained by the presence of multiple TLSs, complicating the calibration of large-scale processors [11].
Various techniques have been developed to address these challenges [12, 13, 14, 15]. The Derivative Removal by Adiabatic Gate (DRAG) method [16, 17, 18], in particular, was a pulse shaping technique designed to suppress leakage transitions to higher energy states during single-qubit operations. By adding the derivative of the original pulse envelope to the quadrature component, DRAG removes unwanted diabatic leakage transitions, which, in the weak drive regime, show up as a spectral hole at a frequency determined by the prefactor of the derivative. The versatility of DRAG has been investigated in a variety of scenarios, including three-level lambda systems [19], cross-resonance gates [20, 21, 22], and frequency-crowded multilevel systems [23]. However, traditional DRAG corrections face challenges when dealing with weakly detuned transitions, as they struggle to effectively remove spectral components close to the target transition [21].
In this work, we propose and experimentally demonstrate a robust approach to mitigating undesired transitions caused by quantum crosstalk during single-qubit operations. Unlike conventional DRAG techniques, our method employs the dual-DRAG protocol, which creates symmetric spectral holes around the target transition frequency. This approach significantly reduces off-resonance effects during pulses and enables the pulse calibration to suppress unwanted transitions that are only slightly detuned. Residual off-resonance effects are further corrected using compensatory virtual-Z (VZ) gates. Experimental results validate the technique, demonstrating an order-of-magnitude suppression of crosstalk-induced excitations between two superconducting qubits detuned by approximately 40 MHz with a 25-ns pulse across various coupling strengths. Furthermore, we show its effectiveness in reducing excitations from a strongly coupled TLS and enhancing single-qubit gate fidelity across a broad frequency range, extending down to 20 MHz.
Figure 1: (a) Schematic diagram illustrating the unintended coupling ($g$) that can occur between two qubits or between a qubit and a spurious TLS. (b) Energy level diagram for the combined system of $\rm Q_{0}$ and $\rm Q_{1}$ (or TLS). Due to state dressing, driving $\rm Q_{0}$ with an amplitude $\Omega$ also cross-drives $\rm Q_{1}$ with a reduced amplitude of ±$\pm g\Omega/\Delta_{0}$ in the weak coupling limit ($g\ll|\Delta_{0}|$). Cross-driving can induce unwanted transitions in $\rm Q_{1}$ when using a pulse with finite width. (c) Time-domain pulse profiles and (d) their normalized Fourier spectra for a 25-ns pulse, showing its original sine-to-the-fourth-power form ($\rm\Omega^{(0)}$), the first-order DRAG-corrected pulse ($\rm\Omega^{(1)}$), and the spectrally balanced pulse with dual-DRAG corrections ($\rm\Omega^{(2)}$). Note that, in the time domain, the real part of $\Omega^{(0)}$ overlaps with that of $\Omega^{(1)}$, while the imaginary part of $\Omega^{(0)}$ overlaps with that of $\Omega^{(2)}$. The vertical dashed lines in (d) indicate the corresponding peak positions. In this case, the frequency to block is assumed to be $\rm\Delta/2\pi=40\ MHz$.
Quantum crosstalk — Parasitic couplings are a common challenge in practical quantum hardware, often arising between qubits that are not intentionally connected. For example, in solid-state devices like superconducting quantum processors, qubits are typically designed for nearest-neighbor connections. However, the long-range nature of electromagnetic interactions can lead to unintended couplings between physically distant qubits, as shown in Fig. 1(a). These interactions can also occur through shared modes, such as chip or box modes. Additionally, unwanted couplings can exist between qubits and spurious quantum systems in the environment, such as TLSs. These uncontrollable systems present a major obstacle in the development and calibration of state-of-the-art quantum hardware.
Consider a parasitic exchange-type coupling $g$ between $\rm Q_{0}$ and $\rm Q_{1}$ (or a TLS). Due to the hybridization of the $\left|\rm ge\right\rangle$ and $\left|\rm eg\right\rangle$ states, driving one qubit inevitably induces a cross-driving effect on the other qubit or TLS, but with a reduced amplitude of $g\Omega/\Delta_{0}$, where $\Omega$ is the original drive amplitude applied to $\rm Q_{0}$, and $\Delta_{0}=\omega_{\rm ge}-\omega_{\rm eg}$ is the detuning between the unwanted and target transitions, as illustrated in Fig. 1(b). Natural spectral broadening caused by finite pulse widths can lead to unwanted transitions, such as $\left|\rm gg\right\rangle\leftrightarrow\left|\rm ge\right\rangle$, which degrade the performance of quantum operations. This crosstalk phenomenon is inherently quantum mechanical; the crosstalk Hamiltonian is of the ZX type, exhibiting opposite signs depending on the state of $\rm Q_{0}$. This contrasts with classical signal crosstalk, where the Hamiltonian is of the IX type. As a result, quantum crosstalk cannot be simply corrected by applying a cancellation signal [24].
Interestingly, this quantum crosstalk effect can also be harnessed for beneficial purposes. For instance, the cross-resonance effect, which is achieved by intentionally driving at the resonant frequency of $\rm Q_{1}$, is employed in generating a type of two-qubit entangling gate, the cross-resonance gate [25, 26, 27].
Dual-DRAG — A straightforward solution to suppress an unwanted transition is to remove the spectral components at that transition frequency. A famous example is the DRAG method, which was initially proposed to address leakage transitions to higher energy states in weakly anharmonic systems, such as transmon qubits.
In the DRAG framework, a first-order derivative of the original pulse envelope (which can be complex) is added as a quadrature component. The amplitude factor of the derivative pulse is expressed as $1/\Delta$, where $\Delta$ is the DRAG parameter in units of angular frequency. An example is illustrated in Fig. 1(c). From a frequency-domain perspective, this DRAG correction creates a spectral hole at a frequency $\Delta$ relative to the drive frequency $\omega_{0}$, as depicted in Fig. 1(d). When $\Delta$ is set equal to the qubit anharmonicity $\alpha$, it effectively reduces spectral components near the qubit’s $\left|\rm e\right\rangle$-$\left|\rm f\right\rangle$ transition frequency—typically 200-300 MHz below the $\left|\rm g\right\rangle$-$\left|\rm e\right\rangle$ transition for transmon qubits—thereby suppressing leakage transitions. This technique has also been applied in cross-resonance gates to mitigate off-resonance errors [21] or parasitic transitions [22].
Avoiding weakly detuned transitions during single-qubit operations, which are only a few tens of MHz away, has proven to be challenging [21]. For instance, consider a 25-ns pulse, as shown in Fig. 1(c,d). While a single DRAG correction generates a spectral hole at the frequency defined by $\Delta$, it also shifts the spectral weight to the opposite side. As shown by the example in Fig. 1(d), it leads to an average drive frequency that is about 30 MHz lower than the original drive frequency $\omega_{0}$ given the frequency to block at $\Delta/2\pi=40$ MHz. This effect may be partially mitigated by introducing a constant drive frequency detuning $\eta$ during the pulse [18], i.e. modulating the nominal drive frequency from $\omega_{0}$ to $\omega_{0}+\eta$. However, when the DRAG-induced off-resonance effect is this large, the perturbative treatment in the DRAG theory breaks down. Often, calibrating a high-fidelity gate under such strong drive-induced detuning is experimentally infeasible.
To address this issue, we adopt the recursive DRAG scheme, where DRAG corrections are applied sequentially according to
$\Omega^{(n)}=\Omega^{(n-1)}-i\frac{\dot{\Omega}^{(n-1)}}{\Delta_{n}}\;,$ (1)
using a set of DRAG parameters $\{\Delta_{1},\Delta_{2},\dots,\Delta_{n}\}$ [28, 22]. Here, $\Omega^{(0)}$ represents the original pulse shape without DRAG. The final pulse $\Omega^{(n)}$ generates multiple spectral holes at the specified frequencies. By two successive DRAG applications at ±$\pm\Delta$, we obtain a pulse envelope that involves a second derivative,
$\Omega^{(2)}=\Omega^{(0)}+\frac{\ddot{\Omega}^{(0)}}{\Delta^{2}}\;.$ (2)
As shown in Fig. 1(d), the spectrum $\hat{\Omega}^{(2)}(\omega)$ remains centered at $\omega_{0}$, while spectral components at $\omega_{0}\pm\Delta$ are eliminated. This dual-DRAG application at mirrored frequencies suppresses weakly detuned transitions and significantly reduces drive-induced off-resonance (see the Supplementary Materials [29] for additional details). For the example presented, we choose the original pulse shape to be a sine raised to the fourth power. This ensures that the first three derivatives vanish at the pulse’s beginning and end, thereby avoiding sharp edges or the need for truncation.
Figure 2: (a) Calibrating the pulse amplitude of the $\rm R$ gate ($t_{g}=25$ ns) which drives the qubit from its ground state to an equator state, as indicated by the arrow. The drive detuning is set at $\eta/2\pi=23\,{\rm MHz}$. The inset plots are the circuit and a sketch of the state evolution on the Bloch sphere. (b) Calibrating the compensating virtual-Z phase after the $\rm R$ gate, which yields an $\rm\sqrt{X}$ gate. By applying the combination twice, the qubit is rotated from the ground state to the excited state, as indicated by the arrow. The inset is the circuit. (c) Experimental results (top panel) showing the calibrated drive amplitude and virtual-Z phase for different drive detunings, compared with numerical simulations (bottom panel). (d) Measured error per Clifford (EPC) from randomized benchmarking using pulses with virtual-Z compensation (blue) for different drive detunings. Results using detuning compensation (without the appended virtual-Z phase, shown in red) are also included for comparison.
Correcting off-resonance error using virtual-Z compensation — Although the dual-DRAG method substantially reduces the deviation of spectral weighting, system nonidealities such as the presence of higher energy levels and pulse distortions can lead to residual off-resonance effects. To address this, we adopt a post-correction approach that compensates for the overall effect of an off-resonant pulse by appending a VZ gate. We validate this protocol on a device similar to that described in Ref. [30] (see Supplementary Materials [29] for more information). For all experiments in this work unless specified otherwise, we constantly apply one DRAG correction at the qubit anharmonicity $\alpha/2\pi\approx-190$ MHz to prohibit leakage transition to the $\left|\rm f\right\rangle$ state.
Throughout our work, we employ the U3 decomposition rule to synthesize arbitrary single-qubit operations using two $\rm\sqrt{X}$ gates interspersed with VZ gates [31]. Consequently, it suffices to calibrate only the $\rm\sqrt{X}$ gate. It can be shown that any rotation $\rm R$, which maps the qubit from its ground state to the equator (XY plane) of the Bloch sphere, is equivalent to a $\frac{\pi}{2}$ rotation around an axis in the XY plane when combined with a virtual Z or phase rotation (see Supplementary Materials [29] and Ref. [31]). Importantly, since single-qubit gates depend only on relative phase, a $\frac{\pi}{2}$ gate of arbitrary phase can serve as $\rm\sqrt{X}$. Thus, calibrating such an $\rm R$ operation enables the synthesis of any single-qubit gate with the aid of VZ gates.
To calibrate the $\rm R$ gate for a given pulse shape, duration, carrier frequency, and DRAG parameters, we scan the drive amplitude until the excited-state population of the qubit (initialized in its ground state) reaches 0.5 (Fig. 2(a)). By concatenating two $\rm R$ gates with interspersed VZ gates, we then determine the VZ phases required to achieve maximal excitation Fig. 2(b). Based on these initial values, we then fine-tune these parameters using standard pulse train techniques (see supplementary Materials [29] for details). Figure 2(c) shows experimental results for the calibrated parameters $\Omega$ (drive amplitude) and $\phi$ (VZ phase) under intentional drive detunings, which align closely with numerical simulations. Using these VZ-compensated $\rm\sqrt{X}$ gates, we perform randomized benchmarking (RB) [32]. The extracted errors per Clifford (EPC), shown in Fig. 2(d), demonstrate consistent performance across a broad detuning range of over ±$\pm 20$ MHz. Compared to detuning compensation during pulses, the VZ compensation approach is better suited to handle more general and complex scenarios. In practice, when off-resonance effects are strong, balancing the correction of off-resonance errors while avoiding spurious transitions becomes challenging. Our method addresses this challenge effectively.
Figure 3: (a) Experimental setup consisting of two transmon qubits and a tunable coupler, which is also a transmon qubit. The coupler is used to modulate the effective coupling strength between the qubits by adjusting its frequency. (b) Pulse sequence designed for detecting spurious transitions in a nearby qubit during single-qubit gate operations. Repeated $\pi$ gates, each consisting of two $\pi/2$ pulses and spaced by a waiting time $\tau$, are applied to $\rm Q_{0}$. (c) Measured excited-state populations of $\rm Q_{0}$ (top panel) and $\rm Q_{1}$ (bottom panel) using the detection sequence, as a function of the waiting time $\tau$. Results are shown with and without dual-DRAG. Both cases include DRAG correction at the leakage transition to the second excited state $\left|2\right\rangle$. The number of $\pi$ gate pairs is $N=50$. Here, $g/2\pi=1.0~{}\rm{MHz}$, $\Delta_{0}/2\pi=45~{}\rm{MHz}$, and $t_{g}=25~{}\rm{ns}$. (d) Randomized benchmarking results for $\rm Q_{0}$, comparing cases with and without dual-DRAG. (e) Simultaneously monitored excited-state populations of $\rm Q_{1}$. The excitation rates per Clifford gate (ExPC) are $8.4\times 10^{-6}$ with dual-DRAG and $1.6\times 10^{-4}$ without dual-DRAG. (f) Error per Clifford for different coupling strengths, with a fixed detuning of $\Delta_{0}/2\pi=45~{}\rm{MHz}$. (g) Excitation per Clifford for varying coupling strengths. The solid lines are the simulated ExPC averaged over 24 single-qubit Cliffords implemented with the U3 decomposition. The excitation rates are proportional to $g^{2}$ (see Supplementary Materials [29] for details).
Suppressing spurious transitions between qubits — Next, we validate the effectiveness of the dual-DRAG method in suppressing unwanted transitions due to quantum crosstalk. Our investigation is performed on two transmon qubits on the same device. Between them is a tunable coupler—another transmon qubit—which adjusts the effective qubit-qubit coupling strength $g$ via the coupler frequency $\omega_{\rm c}$ (Fig. 3(a)). The qubit frequencies are tunable; their detuning is fixed at $\Delta_{0}/2\pi=45~{}\text{MHz}$. At first, we set $g/2\pi=1$ MHz.
For efficient detection of weak transitions ($g\Omega/\Delta_{0}<\Delta_{0}$), we employ a similar pulse sequence adopted for calibrating the classical microwave crosstalk [33] (Fig. 3(b)). An even number ($2N$) of $\rm X$ gates are applied to the control qubit $\rm Q_{0}$, keeping it in the ground state. Due to non-zero coupling, the spectator qubit $\rm Q_{1}$ is coherently excited by these pulses. A uniform waiting time $\tau$ is added after each $\pi$ gate, during which the wavefunction amplitude of $\rm Q_{1}$’s excited state (e.g., $\left|\rm ge\right\rangle$) accumulates a relative phase $e^{-\mathrm{i}\Delta_{0}\tau}$. At specific $\tau$, the transition amplitudes interfere constructively, amplifying $\rm Q_{1}$’s excitation.
Figure 3(c) shows the measured total populations of all excited states ($P^{\rm e}_{i}$, $i=0,1$), including $\left|\rm e\right\rangle$ and $\left|\rm f\right\rangle$, for both qubits as a function of $\tau$. Without dual-DRAG correction, strong periodic peaks appear on both qubits at intervals of $\sim 22.2~{}\text{ns}$, matching $2\pi/\Delta_{0}$ and confirming our expectation. Since $\rm X$ gates are applied in pairs, transitions to the excited state of $\rm Q_{1}$ — whether through $\left|\rm gg\right\rangle\to\left|\rm ge\right\rangle$ or $\left|\rm eg\right\rangle\to\left|\rm ee\right\rangle$ — result in the $\left|\rm ee\right\rangle$ state in the end. Additionally, smaller secondary peaks on $\rm Q_{1}$, halfway between primary peaks and sharing the same periodicity, are attributed to residual $\rm IX$ interaction arising from both quantum and classical crosstalk [29].
We repeat measurements with dual-DRAG correction at ±$\pm\Delta$. The parameter $\Delta$ is initially set to $\Delta_{0}$ and fine-tuned by minimizing excitation peaks. Using three DRAG parameters $\{\alpha,\Delta,-\Delta\}$, all prominent excitation peaks vanish (Fig. 3(c)), demonstrating effective suppression. Randomized benchmarking on $\rm Q_{0}$ (Fig. 3(d,e)) shows a slight error reduction (error per Clifford: $1.60\times 10^{-3}\to 1.42\times 10^{-3}$). Simultaneously, $\rm Q_{1}$’s excitation rate per Clifford (ExPC) drops by an order of magnitude from $(1.6\pm 0.2)\times 10^{-4}$ to $(0.8\pm 0.3)\times 10^{-5}$. Although the effect in EPC does not seem to be significant, the induced correlated excitations, in particular between distant qubits, complicate the error model and may severely affect error correction codes.
We further test the method across coupling strengths using the tunable coupler. The improvement in gate error rates using dual-DRAG becomes more pronounced at larger $g$ (Fig. 3(f)). This trend aligns with the measured spurious excitation rates, which scale quadratically with $g$ as expected from the cross-driving amplitude that scales with $g/\Delta_{0}$. At large $g$, these spurious excitations dominate the gate errors, yet dual-DRAG consistently suppresses them by an order of magnitude across the whole range (Fig. 3(g)). These results demonstrate the effectiveness of spectrally balanced DRAG in suppressing weakly detuned transitions and enhancing gate fidelities.
Figure 4: (a) Spectroscopy of a tunable transmon qubit coupled to a spurious TLS, showing an avoided crossing of approximately 10 MHz. The horizontal axis represents the voltage of the flux bias pulse applied to the coupler qubit. (b) Error per Clifford of the qubit as a function of detuning between the qubit and the TLS. Here $t_{g}=25$ ns. (c) Excitation per Clifford of the TLS for different gate times at $\Delta_{0}/2\pi=42~{}\rm{MHz}$. The solid lines are twice the simulated $\rm\sqrt{X}$ gate errors. (d) Excitation per Clifford of the TLS for varying detunings. The solid lines are twice the simulated $\rm\sqrt{X}$ gate errors.
Applications to the qubit-TLS system — We extend our method to suppress spurious transitions in systems where $\rm Q_{0}$ is coupled to a parasitic TLS. Such TLSs with long relaxation times are particularly problematic, as their unintended excitations can accumulate over time and persistently degrade subsequent gate operations. The specific TLS studied here exhibits a strong coupling ($g/2\pi=5\,\text{MHz}$) to a transmon qubit, as evidenced by the avoided crossing in the measured qubit spectrum shown in Fig. 4(a).
To characterize the dependence on detuning of our method, we calibrate the $\rm\sqrt{X}$ gate at varying qubit-TLS detunings $\Delta_{0}$ using the same protocol developed for coupled qubits. Figure 4(b) shows RB results comparing dual-DRAG corrections (applied at ±$\pm\Delta$) to uncorrected pulses across $20\,\text{MHz}\leq\Delta_{0}/2\pi\leq 85\,\text{MHz}$. For $\Delta_{0}/2\pi<20\,\text{MHz}$, spectral overlap between the drive frequency and the TLS transition hinders the convergence of parameters during calibration, and the dual-DRAG correction overcuts the spectral components around the target frequency, making the drive ineffective. At $\Delta_{0}/2\pi>85\,\text{MHz}$, TLS-induced gate errors diminish to a negligible level.
We further quantify TLS excitation rates across gate time $t_{g}$ and detuning $\Delta_{0}$ (Fig. 4(c,d)). To measure TLS excitations, we: (1) reset the qubit to its ground state post-RB, (2) apply an iSWAP gate to transfer TLS excitations to the qubit, and (3) perform qubit readout. In $20\,\text{ns}\leq t_{g}\leq 50\,\text{ns}$, dual-DRAG reduces TLS excitation rates by an order of magnitude compared to uncorrected pulses. Numerical simulations of single-pulse dynamics reproduce the observed decreasing trend, though deviations may arise from interference between pulses and spectral smoothing due to the presence of dephasing noise.
The measured excitation rates for $20\,\text{MHz}\leq\Delta_{0}/2\pi\leq 85\,\text{MHz}$ show a comparable improvement from the use of dual-DRAG. The rapidly enhanced suppression with increasing $\Delta_{0}$ results from the diminishing spectral weight at larger $\Delta_{0}$, and the reduced cross-drive amplitude that scales with $g/\Delta_{0}$.
Discussion — We have demonstrated a spectrally balanced pulse-shaping technique that suppresses spurious transitions in superconducting qubits. Our method reduces crosstalk errors by an order of magnitude and mitigates single-qubit gate errors from strongly coupled TLS. By engineering symmetric spectral holes, it effectively suppresses weakly detuned transitions by mitigating the off-resonance effect during pulses. The technique integrates seamlessly with virtual-Z gates, further improving gate fidelity.
This approach directly addresses key scalability challenges in a frequency-crowded quantum processor: parasitic qubit couplings and TLS-induced errors, both of which limit frequency allocation. By mitigating these effects, our work can enhance frequency planning flexibility and provide a framework for modeling microwave gate errors in complex environments. The hardware-agnostic nature of the dual-DRAG protocol makes it broadly applicable to other quantum architectures, including quantum dots, NV centers, and trapped ions, positioning it as a versatile tool for high-fidelity quantum control.
Note added. — During the preparation of this manuscript, we became aware of a concurrent work [34] that independently develops a related pulse-shaping technique using symmetrically filtered spectra to address classical crosstalk.
Acknowledgements.
The authors would like to thank Yasunobu Nakamura and Peng Zhao for fruitful discussions. This work was supported by the National Natural Science Foundation of China (Grants No. 12404558, No. 12322413, No. 92365206, and No. 92476206), Innovation Program for Quantum Science and Technology (Grants No.2021ZD0301802).
Supplementary Materials for “Suppressing spurious transitions using spectrally balanced pulse”
Supplementary Materials for “Suppressing spurious transitions using spectrally balanced pulse”
I Experimental setup
The superconducting quantum processor used in this study and the experimental setup closely mirror those described in Ref. [30]. Qubits $\rm Q_{0}$ and $\rm Q_{1}$ are frequency-tunable transmon qubits, coupled through a tunable coupler. The coupling strength between the qubits can be adjusted by tuning the coupler’s frequency. Qubit $\rm Q_{0}$ is also employed in the TLS experiment and is coupled to the relevant TLS. Table 1 lists the maximum ($\omega^{\rm max}$) and typical operating ($\omega^{\rm idle}$) frequencies of the qubits and the TLS $\left|0\right\rangle\rightarrow\left|1\right\rangle$ transition, as well as the anharmonicities of the qubits. Additionally, it provides the $T_{1}$ relaxation times measured at both maximum and operating frequencies, and the $T_{2}$ times obtained via Ramsey interferometry at those frequencies.
$\rm Q_{0}$ $\rm Q_{1}$ TLS
  $\omega^{\rm max}/2\pi$ (GHz)   3.91   3.76  3.84
  $\omega^{\rm idle}/2\pi$ (GHz)   3.76   3.72  3.84
  $\alpha/2\pi$ (MHz)   -194.6   -193.2 --
  $T_{1}^{\rm max}$ ($\mu$s)   78.6   72.6 --
  $T_{1}^{\rm idle}$ ($\mu$s)   78.1   76.8  14.32
  $T_{2R}^{\rm max}$ ($\mu$s)   37.3   22.2 --
  $T_{2R}^{\rm idle}$ ($\mu$s)   1.1   5.5 --
Table 1: Parameters describing the two qubits and qubit-TLS system studied in our experiment.
II Theoretical model
The Hamiltonian for the two-qubit system is ($\hbar=1$)
$\begin{split}H=&\ H_{0}+H_{D},\\ H_{0}=&\ \sum_{i=0,1}(\omega_{i}a_{i}^{\dagger}a_{i}+\frac{\alpha_{i}}{2}a_{i}^{\dagger}a_{i}^{\dagger}a_{i}a_{i})+g(a_{0}^{\dagger}a_{1}+H.c.),\\ H_{D}=&\ \Omega(t)(a_{0}^{\dagger}+a_{0}).\end{split}$ (3)
$H_{0}$ is the drift Hamiltonian, with a constant coupling strength $g$ between the two qubits, while $H_{D}$ is the driving Hamiltonian, representing the single-qubit operation on $\rm Q_{0}$. In the absence of a driving pulse, the dressed states of the Hamiltonian $H_{0}$ are denoted as $|{\rm ij}\rangle$, where the eigenfrequencies are given by $\Lambda_{{\rm ij}}$ ($\rm i,j=g,e$). The driving frequency applied to $\rm Q_{0}$ is set to $\Lambda_{\rm eg}$. In the dressed state basis ${|{\rm gg}\rangle,|{\rm ge}\rangle,|{\rm eg}\rangle,|{\rm ee}\rangle}$, and using the rotating wave approximation along with the first-order approximation, we have
$H_{r}=\left[\begin{array}[]{cccc}0&\frac{\Omega g}{2\Delta_{0}}&\frac{\Omega}{2}&0\\ \frac{\Omega g}{2\Delta_{0}}&\delta+\Delta_{0}&0&\frac{\Omega}{2}\\ \frac{\Omega}{2}&0&\delta&-\frac{\Omega g}{2\Delta_{0}}\\ 0&\frac{\Omega}{2}&-\frac{\Omega g}{2\Delta_{0}}&2\delta+\Delta_{0}\end{array}\right],$ (8)
where $\Delta_{0}=\Lambda_{\rm ge}-\Lambda_{\rm eg}$, $\delta=\Lambda_{\rm eg}-\omega_{{\rm d}}$ and $\omega_{{\rm d}}$ is the driving frequency.
For simplicity, in the Hamiltonian model and DRAG correction analysis (Section V), we only consider the ideal two-level system. But for the transmon qubits, impacted by the higher energy levels, the Hamiltonian of the quantum crosstalk has $\rm ZX$ and $\rm IX$ components [27]. So that, in Section IV, we give the analysis for both $\rm ZX$ and $\rm IX$ interactions.
III Calibration of $\rm\sqrt{X}$ gate with VZ compensation
III.1 Demonstration of the decomposition of a $\frac{\pi}{2}$ gate
The single-qubit operation that rotates an initial state $\left|\rm g\right\rangle$ to the equator of the Bloch sphere can be expressed as:
$\begin{split}R&=\left[\begin{array}[]{cc}\cos{\frac{\pi}{4}}&-ie^{i\lambda}\sin{\frac{\pi}{4}}\\ -ie^{i\theta}\sin{\frac{\pi}{4}}&e^{i(\lambda+\theta)}\cos{\frac{\pi}{4}}\end{array}\right].\end{split}$ (9)
A rotation gate around the Z-axis with a rotation angle of $\lambda+\theta$ can be expressed as:
$\begin{split}R_{z}&=\left[\begin{array}[]{cc}1&0\\ 0&e^{-i(\lambda+\theta)}\end{array}\right].\end{split}$ (10)
Applying $R$ and $R_{z}$ in sequence, we obtain:
$\begin{split}R_{z}R&=\left[\begin{array}[]{cc}\cos{\frac{\pi}{4}}&-ie^{i\lambda}\sin{\frac{\pi}{4}}\\ -ie^{-i\lambda}\sin{\frac{\pi}{4}}&\cos{\frac{\pi}{4}}\end{array}\right]\\ &=\frac{1}{\sqrt{2}}\left[\begin{array}[]{cc}1&-ie^{i\lambda}\\ -ie^{-i\lambda}&1\end{array}\right]\end{split}$ (11)
The matrix representation of the operation $e^{-i\frac{\pi}{4}(\cos{\Phi}\sigma_{x}+\sin{\Phi}\sigma_{y})}$ is
$\begin{split}R_{\Phi}(\frac{\pi}{2})&=\frac{1}{\sqrt{2}}\left[\begin{array}[]{cc}1&-ie^{-i\Phi}\\ -ie^{i\Phi}&1\end{array}\right].\end{split}$ (12)
The equations (11) and (12) are equivalent as long as we set $\lambda=\Phi$. Thus, the arbitrary single-qubit operation $e^{-i\frac{\pi}{4}(\cos{\Phi}\sigma_{x}+\sin{\Phi}\sigma_{y})}$, which we denote as $R_{\Phi}(\frac{\pi}{2})$, can be decomposed into two operations: $R$ and $R_{z}$. In this case, $R_{\Phi}(\frac{\pi}{2})$ is simply a $\frac{\pi}{2}$ gate.
III.2 General flow for the calibration of $\rm\sqrt{X}$ gate
In the experiment, the $\rm\sqrt{X}$ gate is calibrated using the following processes.
[…]
Take $\Omega_{0}$ and $\phi_{0}$ as the initial driving amplitude and VZ compensation. Repeat steps 3 and 4 with increasing values of $n$ until reaching the maximum value of $n$ before decoherence occurs.
3. Prepare the initial state in $\left|\rm g\right\rangle$, apply $24n+$ 2 gates to be calibrated, and scan the driving amplitude. Take the value of the driving amplitude that corresponds to the maximum of $P^{\rm e}$ as the updated $\Omega$.
4. Prepare the initial state in $\left|\rm g\right\rangle$, apply 2 gates to be calibrated, followed by 2 conjugate gates of the gate to be calibrated, and repeat this process $n$ times. Scan the value of $\phi$ and take the value that corresponds to the maximum of $P^{\rm g}$ as the updated $\phi$.
III.3 Calibration of $\rm\sqrt{X}$ gate with dual-DRAG
We calibrate the $\rm\sqrt{X}$ gate using dual-DRAG to achieve optimal suppression performance. First, we choose typical initial values for the driving amplitude $\Omega$, VZ compensation $\phi$, and the DRAG parameters $\{\alpha,\Delta,-\Delta\}$.
Second, we scan the DRAG point $\Delta$ to find the optimal value using two calibration circuits. The first circuit is the pulse train shown in Fig. 3(b) of the main text. By measuring the spurious excitations around one of the strong peaks, we scan the $\Delta$ value around the predicted value, $\Delta_{0}$. The minimum value of $P^{\rm e}_{1}$ corresponds to the optimal $\Delta$, as shown in Fig. 5. We also apply the second circuit in another case. In this case, we implement the RB circuit on qubit $\rm Q_{0}$ with 300 cycles, and measure the population $P^{\rm e}$ of $\rm Q_{1}$ at the end of the RB sequence (for qubit-TLS system, we will measure the qubit after the qubit reset gate and qubit-TLS iSWAP gate). By scanning the value of $\Delta$, the minimum measured population corresponds to the optimal value of $\Delta$, another case using this calibration method is shown in Fig. 6.
Third, using the optimal DRAG set $\{\alpha,\Delta,-\Delta\}$, we calibrate the $\rm\sqrt{X}$ gate following the calibration flow outlined in Sec. III.2.
Figure 5: Calibration processes with the pulse train method. Measured $P^{\rm e}$ of $\rm Q_{1}$ using the same detection circuit as in Fig. 3(b) of the main text, but with dual-DRAG, plotted against the waiting time $\tau$ and the spectral hole-related detuning $\Delta$. The white dashed line indicates the optimized $\Delta$ value used for subsequent experiments.
Figure 6: Calibration processes with the RB circuit. Measured $P^{\rm e}$ of the qubit using dual-DRAG with the RB sequence, setting the number of Cliffords to 300, plotted against the spectral hole-related detuning $\Delta$. The black dashed line indicates the optimized $\Delta$ value used for subsequent experiments.
IV Explanations for the detected peaks
In the experiments for the detection of the spurious transitions, we observed two different groups of peaks with different locations and heights. In this section, we would like to give the detailed analysis for the mechanism of the peaks.
IV.1 ZX crosstalk
The unitary matrix for the X operation on the left qubit with ZX crosstalk can be written as
$U_{X}^{(ZX)}$$\approx$ $\left[\begin{array}[]{cccc}0&-iA&-i&iB\\ -iA&0&-iB&-iC\\ -i&-iB&0&iA\\ iB&-iC&iA&0\end{array}\right]$ (17)
with the approximation of $|g/\Delta_{0}|\ll 1$, where $A=\frac{\Omega^{2}g}{-2\Delta_{0}(\Delta_{0}^{2}-\Omega^{2})}+\frac{\Omega^{2}g}{-2\Delta_{0}(\Delta_{0}^{2}-\Omega^{2})}e^{-i\Delta_{0}T_{g}}$, $B=\frac{\Omega g}{-2(\Delta_{0}^{2}-\Omega^{2})}+\frac{\Omega g}{-2(\Delta_{0}^{2}-\Omega^{2})}e^{-i\Delta_{0}T_{g}}$, $C=e^{-i\Delta_{0}T_{g}}$, $T_{g}$ is the gate time. Here we assume that, the driving pulse is a square pulse with a real and constant amplitude $\Omega$ and a driving frequency equal to $\Lambda_{\rm eg}$. We dose not distinguish the difference of the frequency detuning between the bare states $\omega_{\rm e}-\omega_{\rm g}$ and dressed states $\Lambda_{\rm ge}-\Lambda_{\rm eg}$. The unitary matrix for the buffer process with the buffer time $T_{b}$ can be written as
$U_{bf}\approx\left[\begin{array}[]{cccc}1&0&0&0\\ 0&e^{-i\Delta_{0}T_{b}}&0&0\\ 0&0&1&0\\ 0&0&0&e^{-i\Delta_{0}T_{b}}\end{array}\right]$ (22)
with the approximation of $|g/\Delta_{0}|\ll 1$.
Then we can calculate the unitary matrix elements for the sequence of the operations as with the approximation of $|A|,|B|\ll 1$
$\begin{split}|[U_{bf}U_{X}^{(ZX)}U_{bf}U_{X}^{(ZX)}]_{[{\rm gg}->{\rm ge}]}^{n(n\geq 1)}|&\approx|-Be^{-i\Delta_{0}T_{b}}[1-e^{-i\Delta_{0}(T_{g}+T_{b})}][1+e^{-i2\Delta_{0}(T_{g}+T_{b})}+e^{-i4\Delta_{0}(T_{g}+T_{b})}+...]|\\ &=|B^{\prime}e^{-i\Delta_{0}T_{b}}(1+e^{-i\Delta_{0}T_{g}})(1-e^{-i\theta})||\frac{\sin{n\theta}}{\sin{\theta}}|,\\ |[U_{bf}U_{X}^{(ZX)}U_{bf}U_{X}^{(ZX)}]_{[{\rm gg}->{\rm ee}]}^{n(n\geq 1)}|&\approx|-Ae^{-i\Delta_{0}T_{b}}[1-e^{-i\Delta_{0}(T_{b}+T_{g})}][1+e^{-i2\Delta_{0}(T_{g}+T_{b})}+e^{-i4\Delta_{0}(T_{g}+T_{b})}+...]|\\ &=|A^{\prime}e^{-i\Delta_{0}T_{b}}(1+e^{-i\Delta_{0}T_{g}})(1-e^{-i\theta})||\frac{\sin{n\theta}}{\sin{\theta}}|,\end{split}$ (23)
where $A^{\prime}=\frac{\Omega^{2}g}{-2\Delta_{0}(\Delta_{0}^{2}-\Omega^{2})}$, $B^{\prime}=\frac{\Omega g}{-2(\Delta_{0}^{2}-\Omega^{2})}$, $\theta=\Delta_{0}(T_{g}+T_{b})$. The results show that, when $\theta=m\pi$ ($m\in N$), the term of $|\frac{\sin{n\theta}}{\sin{\theta}}|$ is in the peak value. However, when $\theta=2k\pi$ ($k\in N$), there is $01-e^{-i\theta}=$ 0, then only when $\theta=(2k+1)\pi$, we can detect the peaks in the states $\left|\rm ge\right\rangle$ and $\left|\rm ee\right\rangle$. The buffer time between two adjacent peaks is $\Delta T_{b}=|2\pi/\Delta_{0}|$. We can also know that, for our sequence, when the initial state is $\left|\rm gg\right\rangle$, the peaks of $\left|\rm ee\right\rangle$ represents the interaction of the first-order process between $\left|\rm gg\right\rangle$ and $\left|\rm ge\right\rangle$ or $\left|\rm eg\right\rangle$ and $\left|\rm ee\right\rangle$, while the peaks of $\left|\rm ge\right\rangle$ indicates the second-order interaction between $\left|\rm gg\right\rangle$ and $\left|\rm ee\right\rangle$ or $\left|\rm ge\right\rangle$ and $\left|\rm eg\right\rangle$.
IV.2 IX crosstalk
The unitary matrix for the X operation on the left qubit with IX can be written as
$U_{X}^{(IX)}$ $=$ $\left[\begin{array}[]{cccc}0&0&-i&iD\\ 0&0&iD&-iC\\ -i&iD&0&0\\ iD&-iC&0&0\end{array}\right],$ (28)
with the approximation of $|g/\Delta_{0}|\ll 1$ and the assumption of a constant driving amplitude, where $D=\nu(1-e^{-i\Delta_{0}T_{g}})$, $C=e^{-i\Delta_{0}T_{g}}$. $\nu$ is a constant, the value of which depends on the strength for both of the quantum crosstalk induced IX interaction and the microwave crosstalk. with the approximation of $|D|\ll 1$, the calculated unitary matrix elements for the sequence of the detection circuit are
$\begin{split}|[U_{bf}U_{X}^{(IX)}U_{bf}U_{X}^{(IX)}]_{[{\rm gg}->{\rm ge}]}^{n(n\geq 1)}|&\approx|D(1+e^{-i\Delta_{0}(T_{b}+T_{g})})[1+e^{-i2\Delta_{0}(T_{b}+T_{g})}+e^{-i4\Delta_{0}(T_{b}+T_{g})}+...]|\\ &=|\nu(1-e^{-i\Delta_{0}T_{g}})(1+e^{-i\theta})||\frac{\sin{n\theta}}{\sin{\theta}}|\\ |[U_{bf}U_{X}^{(IX)}U_{bf}U_{X}^{(IX)}]_{[{\rm gg}->{\rm ee}]}^{n(n\geq 1)}|&\approx 0\end{split}$ (29)
Compared with the case of ZX interaction, we can get that, the buffer time between two adjacent peaks is also $\Delta T_{b}=|2\pi/\Delta_{0}|$, but only when $\theta=2k\pi$, the peaks in the states $\left|\rm ge\right\rangle$ can be detected. So that, the peaks induced by IX interaction are always resides in the middle of the two peaks induced by the ZX interaction.
V Analysis for DRAG corrections
In this section, we would like to give the analysis for the difference between the traditional DRAG with constant detuning and the spectrally balanced DRAG corrections. For the traditional DRAG, the adiabatic transformation is
$V_{c}=\exp[-iS_{c}(\frac{g}{\Delta_{0}}\sigma_{y}^{\rm gg-ge}+\sigma_{y}^{\rm gg-eg}+\sigma_{y}^{\rm ge-ee}-\frac{g}{\Delta_{0}}\sigma_{y}^{\rm eg-ee})],$ (30)
where $\sigma_{y}^{j-k}=-i|j\rangle\langle k|+i|k\rangle\langle j|$. Then we can get the results of the first order correction with
$\begin{split}S_{c}&=\frac{\Omega^{(0)}}{2\Delta},\\ \Omega^{(1)}&=\Omega^{(0)}-i\frac{\dot{\Omega}^{(0)}}{\Delta}.\end{split}$ (31)
The adiabatic transformation is $H^{V}=VH_{r}V^{\dagger}+i\dot{V}V^{\dagger}$, the effective Hamiltonian with the driving pulse $\Omega^{(1)}$ is
$H^{V}\approx\left[\begin{array}[]{cccc}\frac{\eta[\Omega^{(0)}]^{2}}{4\Delta_{0}^{2}}-\frac{[\Omega^{(0)}]^{2}}{2\Delta_{0}}&-\frac{\eta g\Omega^{(0)}}{2\Delta_{0}^{2}}-\frac{g[\Omega^{(0)}]^{3}}{4\Delta_{0}^{3}}&-\frac{\eta\Omega^{(0)}}{2\Delta_{0}}-\frac{[\Omega^{(0)}]^{3}}{4\Delta_{0}^{2}}+\frac{\Omega^{(0)}}{2}&-\frac{g[\Omega^{(0)}]^{2}}{4\Delta_{0}^{2}}\\ -\frac{\eta g\Omega^{(0)}}{2\Delta_{0}^{2}}-\frac{g[\Omega^{(0)}]^{3}}{4\Delta_{0}^{3}}&\frac{\eta[\Omega^{(0)}]^{2}}{4\Delta_{0}^{2}}+\eta-\frac{[\Omega^{(0)}]^{2}}{2\Delta_{0}}+\Delta_{0}&-\frac{\eta g[\Omega^{(0)}]^{2}}{2\Delta_{0}^{3}}+\frac{3g[\Omega^{(0)}]^{2}}{4\Delta_{0}^{2}}&-\frac{\eta\Omega^{(0)}}{2\Delta_{0}}-\frac{[\Omega^{(0)}]^{3}}{4\Delta_{0}^{2}}+\frac{\Omega^{(0)}}{2}\\ -\frac{\eta\Omega^{(0)}}{2\Delta_{0}}-\frac{[\Omega^{(0)}]^{3}}{4\Delta_{0}^{2}}+\frac{\Omega^{(0)}}{2}&-\frac{\eta g[\Omega^{(0)}]^{2}}{2\Delta_{0}^{3}}+\frac{3g[\Omega^{(0)}]^{2}}{4\Delta_{0}^{2}}&-\frac{\eta[\Omega^{(0)}]^{2}}{4\Delta_{0}^{2}}+\eta+\frac{[\Omega^{(0)}]^{2}}{2\Delta_{0}}&\frac{\eta g\Omega^{(0)}}{2\Delta_{0}^{2}}+\frac{g[\Omega^{(0)}]^{3}}{4\Delta_{0}^{3}}\\ -\frac{g[\Omega^{(0)}]^{2}}{4\Delta_{0}^{2}}&-\frac{\delta\Omega^{(0)}}{2\Delta_{0}}-\frac{[\Omega^{(0)}]^{3}}{4\Delta_{0}^{2}}+\frac{\Omega^{(0)}}{2}&\frac{\eta g\Omega^{(0)}}{2\Delta_{0}^{2}}+\frac{g[\Omega^{(0)}]^{3}}{4\Delta_{0}^{3}}&-\frac{\eta[\Omega^{(0)}]^{2}}{4\Delta_{0}^{2}}+2\eta+\frac{[\Omega^{(0)}]^{2}}{2\Delta_{0}}+\Delta_{0}\end{array}\right].$
Because there is $g/\Delta_{0}\ll 1$, we discard the items with the power of $g/\Delta_{0}$ higher than 1. In our problem, the value of $\Omega^{(0)}$ is not far less than ${\Delta_{0}}$, then the term with $\frac{\Omega^{(0)}}{\Delta_{0}}$ cannot be discarded simply. From $H^{V}[1,1]$ and $H^{V}[3,3]$, we can get the value of the detuning correction $\eta$ as
$\eta=-\frac{[\Omega^{(0)}]^{2}}{\Delta_{0}(1-[\Omega^{(0)}]^{2}/2\Delta_{0}^{2})}$ (33)
According to the experimental and numerical simulation results, the maximum amplitude of the driving pulse is about 30 MHz, then we assume the maximum value of $\Omega^{(0)}/\Delta_{0}\approx 3/4$, the detuning is about $-1.39[\Omega^{(0)}]^{2}/\Delta_{0}$. The item related to the first-order quantum crosstalk is $H^{V}[1,2]=-0.89[\Omega^{(0)}]^{3}g/2\Delta_{0}^{3}$ and $H^{V}[4,3]=-H^{V}[1,2]$, compared with the original item in $H_{r}[1,2]$, there is $H^{V}[1,2]/H_{r}[1,2]=0.89[\Omega^{(0)}]^{2}/\Delta_{0}^{2}\approx 0.5$, which is not far less than 1. That is the reason for the loss of efficacy of the traditional DRAG correction. There is one thing to note, even without considering the phase correction and setting $\eta=0$, the value of $H^{V}[1,2]/H_{r}[1,2]$ about $0.28$, the suppression for the crosstalk induced transitions is still limited.
For the spectrally balanced DRAG corrections, there is
$V_{b}=\exp[-iS_{b1}(\frac{g}{\Delta_{0}}\sigma_{y}^{\rm gg-ge}-\frac{g}{\Delta_{0}}\sigma_{y}^{\rm eg-ee})-iS_{b2}(\frac{g}{\Delta_{0}}\sigma_{x}^{\rm gg-ge}-\frac{g}{\Delta_{0}}\sigma_{x}^{\rm eg-ee})],$ (34)
where $\sigma_{x}^{j-k}=|j\rangle\langle k|+|k\rangle\langle j|$. Then we can get the results of the first and second order corrections with
$\begin{split}S_{b1}&=\frac{\Omega^{(0)}}{2\Delta},\\ S_{b2}&=-\frac{\dot{\Omega}^{(0)}}{2\Delta^{2}},\\ \Omega^{(2)}&=\Omega^{(0)}+\frac{\ddot{\Omega}^{(0)}}{\Delta^{2}}.\end{split}$ (35)
After the adiabatic transformation, there are
$H^{V}\approx\left[\begin{array}[]{cc}A&B\\ C&D\end{array}\right],$ (38)
where
$\begin{split}A&\approx\left[\begin{array}[]{cc}0&-\frac{\eta g\Omega^{(0)}}{2\Delta^{2}}+\frac{i\eta g\dot{\Omega}^{(0)}}{2\Delta^{3}}\\ -\frac{\eta g\Omega^{(0)}}{2\Delta^{2}}-\frac{i\eta g\dot{\Omega}^{(0)}}{2\Delta^{3}}&\Delta+\eta\end{array}\right],\\ B&\approx\left[\begin{array}[]{cc}\frac{\ddot{\Omega}^{(0)}}{2\Delta^{2}}+\frac{\Omega^{(0)}}{2}&\frac{ig\Omega^{(0)}\dot{\Omega}^{(0)}}{2\Delta^{3}}-\frac{g[\Omega^{(0)}]^{2}}{2\Delta^{2}}\\ \frac{ig\Omega^{(0)}\dot{\Omega}^{(0)}}{2\Delta^{3}}+\frac{g[\Omega^{(0)}]^{2}}{2\Delta^{2}}&\frac{\ddot{\Omega}^{(0)}}{2\Delta^{2}}+\frac{\Omega^{(0)}}{2}\end{array}\right],\\ C&\approx\left[\begin{array}[]{cc}\frac{\ddot{\Omega}^{(0)}}{2\Delta^{2}}+\frac{\Omega^{(0)}}{2}&-\frac{ig\Omega^{(0)}\dot{\Omega}^{(0)}}{2\Delta^{3}}+\frac{g[\Omega^{(0)}]^{2}}{2\Delta^{2}}\\ -\frac{ig\Omega^{(0)}\dot{\Omega}^{(0)}}{2\Delta^{3}}-\frac{g[\Omega^{(0)}]^{2}}{2\Delta^{2}}&\frac{\ddot{\Omega}^{(0)}}{2\Delta^{2}}+\frac{\Omega^{(0)}}{2}\end{array}\right],\\ D&\approx\left[\begin{array}[]{cc}\eta&\frac{\eta g\Omega^{(0)}}{2\Delta^{2}}-\frac{i\eta g\dot{\Omega}^{(0)}}{2\Delta^{3}}\\ \frac{\eta g\Omega^{(0)}}{2\Delta^{2}}+\frac{i\eta g\dot{\Omega}^{(0)}}{2\Delta^{3}}&\Delta+2\eta\end{array}\right].\end{split}$ (39)
Here, we also discard the items with the power of $g/\Delta_{0}$ higher than 1. In this case, from $H^{V}[1,1]$ and $H^{V}[3,3]$, we can get the value of detuning with $\eta=0$. So that, without considering the higher order of $g/\Delta_{0}$, the items related to the quantum crosstalk are zero.
VI Fitting function for the excitation rate
From the equations (17) and (28), we can take the exciting operation on the spectator qubit as a small rotation along the $\sigma_{x}$ axis with $0<\gamma_{e}\ll 1$ and
$R_{x}(\gamma_{e})=\left[\begin{array}[]{cc}\sqrt{1-\gamma_{e}}&-i\sqrt{\gamma_{e}}\\ -i\sqrt{\gamma_{e}}&\sqrt{1-\gamma_{e}}\end{array}\right],$ (40)
To extract the excitation rate $\gamma_{e}$, we assume the qubit goes through a series of channels: $\rho_{f}=\prod_{i}O_{i}(\rho_{0})$, where $O_{i}=C_{\Gamma}\bar{C}_{i}$, $C_{\Gamma}$ is an amplitude damping channel with the damping rate $\Gamma$, and $\bar{C}_{i}=\frac{1}{n}\sum_{j=1}^{n}R_{z}(\alpha_{ij})R_{x}(\gamma_{e})R_{z}(\beta_{ij})$. $R_{z}$ is the single-qubit rotating gate around the $\sigma_{z}$ axis with the angle $\alpha$ and $\beta$, which is chosen randomly and when $n\rightarrow+\infty$, there are
[…]
with $\gamma_{e}=\sin^{2}{\theta}$. Then there is
$\begin{split}\bar{C}_{i}(\rho)&=\cos^{2}{\theta}\rho+\frac{1}{2}\sin^{2}{\theta}(\sigma_{x}\rho\sigma_{x}+\sigma_{y}\rho\sigma_{y})\\ &=(1-\gamma_{e})\rho+\frac{1}{2}\gamma_{e}(\sigma_{x}\rho\sigma_{x}+\sigma_{y}\rho\sigma_{y})\end{split}$ (42)
Initially, there is
$\rho_{0}=\left[\begin{array}[]{cc}1&0\\ 0&0\end{array}\right],$ (43)
After the first channel $\bar{C}_{1}$, the desity matrix becomes
$\rho_{0z}=\left[\begin{array}[]{cc}1-\gamma_{e}&0\\ 0&\gamma_{e}\end{array}\right],$ (44)
And then after the amplitude damping channel, the density matrix is
$\rho_{1}=E_{1}\rho_{0z}E_{1}^{\dagger}+E_{2}\rho_{0z}E_{2}^{\dagger},$ (45)
where
$E_{1}=\left[\begin{array}[]{cc}1&0\\ 0&\sqrt{1-\Gamma}\end{array}\right],\,E_{2}=\left[\begin{array}[]{cc}0&\sqrt{\Gamma}\\ 0&0\end{array}\right].$ (46)
Then there is
$\rho_{1}=\left[\begin{array}[]{cc}1-\gamma_{e}+\gamma_{e}\Gamma&0\\ 0&\gamma_{e}(1-\Gamma)\end{array}\right].$ (47)
We assume that, for $\rho_{m-1}$, there is
$\rho_{m-1}=\left[\begin{array}[]{cc}1-p_{1}(m-1)&0\\ 0&p_{1}(m-1)\end{array}\right],$ (48)
where $p_{1}(m-1)$ is the population on the excited state after $m-1$ Clifford gates. We can get the population of the excited state in $\rho_{m}$ as
$\begin{split}p_{1}(m)=&r_{e}-r_{e}\Gamma+(1-\Gamma-2r_{e}+2r_{e}\Gamma)p_{1}(m-1)\\ =&(r_{e}-r_{e}\Gamma)\frac{1-q^{m}}{1-q}\end{split}$ (49)
for $m=1,2,3...$, $p_{1}(0)=0$ and $q=1-\Gamma-2r_{e}+2r_{e}\Gamma$.
VII Explanation of the inter-pulse interference effects
[…]
Figure 7: Simulated excitation rates as a function of gate time for single-qubit operations with the DRAG set $\{\alpha\}$. The green line represents the experimental results. The red line shows twice the excitation rate of the $\rm\sqrt{X}$ gate. The orange line represents the average ExPC from 24 Clifford gates compiled with the U3 scheme. The blue line shows one-third of the excitation rate of averaged single-qubit operations, each consisting of three U3 gates. The average value is computed from 10 random operations.
VIII Comparison of different DRAG sets
To compare the suppression of spurious transitions with different DRAG sets, we present the experimental results for RB and pulse train with four different DRAG sets, as shown in Fig. 8. The ExPC is significantly reduced with the recursive DRAG scheme, and the best suppression occurs with the optimized dual-DRAG, using the DRAG set $\{\alpha,\Delta_{\rm opt},-\Delta_{\rm opt}\}$, as shown in Fig. 8(b). The significant suppression is further confirmed in Fig. 8(c). The irregular lines in the top panel may be attributed to noise in the system.
Figure 8: (a) Randomized benchmarking results for $\rm Q_{0}$, comparing different DRAG sets. (b) Simultaneously monitored excited-state populations of $\rm Q_{1}$. The ExPC values for the DRAG sets $\{\alpha\}$, $\{\alpha,\Delta_{0}\}$, $\{\alpha,\Delta_{0},-\Delta_{0}\}$, and $\{\alpha,\Delta_{\rm opt},-\Delta_{\rm opt}\}$ are $(2.7\pm 0.2)\times 10^{-4}$, $(6.6\pm 1.0)\times 10^{-5}$, $(4.2\pm 1.2)\times 10^{-5}$, and $(3.2\pm 1.2)\times 10^{-5}$, respectively. (c) Measured excited-state populations of $\rm Q_{0}$ (top panel) and $\rm Q_{1}$ (bottom panel), using the detection sequence shown in Fig. 3(b) of the main text, as a function of the waiting time $\tau$. Results are shown for different DRAG sets. All cases include DRAG correction for the leakage transition to the second excited state $\left|2\right\rangle$. The number of $\pi$ gate pairs is $N=50$. Here, $g/2\pi=0.8~{}\rm{MHz}$, $\Delta_{0}/2\pi=42.5~{}\rm{MHz}$, and $t_{g}=24~{}\rm{ns}$.
IX Dependence of ExPC over coupling strength
To verify the functional relationship between the ExPC and the coupling strength $g$, we fit the simulated ExPC shown in Fig. 3(g) of the main text using the fitting function $f(x)=ax^{2}$. The results are shown in Fig. 9. The excitation rates are found to be proportional to $g^{2}$. The fitting parameters are $a_{\rm w/}=2.6\times 10^{-5}$ and $a_{\rm w/o}=3.0\times 10^{-4}$ with and without dual-DRAG, respectively.
Figure 9: (a) Fitting of the ExPC as a function of coupling strength $g$ using the fitting function $f(x)=ax^{2}$. The simulated ExPC values are averaged over 24 single-qubit Cliffords implemented with the U3 decomposition. The results with and without dual-DRAG are represented by blue dots and green squares, respectively, with the corresponding fitting results shown as red and orange solid lines.
X Suppression of the spurious transitions
In Fig. 10, we present the experimental results for TLS excitation rates as a function of gate time ($t_{g}$) with the detuning $\Delta_{0}/2\pi=60~{}{\rm MHz}$, using the circuit shown in Fig. 4(c) of the main text. Similar to the results presented in the main text, we observe that for gate times in the range $20~{}\text{ns}\leq t_{g}\leq 50~{}\text{ns}$, dual-DRAG reduces TLS excitation rates by an order of magnitude compared to uncorrected pulses. Numerical simulations of single-pulse dynamics successfully reproduce the observed decreasing trend.
Figure 10: Excitation per Clifford of the TLS for different gate times at $\Delta_{0}/2\pi=60~{}\rm{MHz}$. The solid lines represent twice the simulated $\rm\sqrt{X}$ gate errors.
XI Leakage suppression with recursive DRAG
Having successfully suppressed spurious transitions using the spectrally balanced DRAG pulse, we now turn our attention to the leakage error on $\rm Q_{0}$ and talk about the suppression of more than one transitions with multiple DRAG corrections. Compared with the conventional DRAG application on the leakage error with the DRAG set of $\{\alpha\}$, when applying the DRAG correction with the set of $\{\Delta,-\Delta\}$, the leakage to the excited states of $\rm Q_{0}$ experiences a sudden increase, as observed by the prominent peaks in FIG. 11(a) with red data line. This increase attributed to the amplification of the driving strength at the frequency of $\omega_{\rm fg}-\omega_{\rm eg}$ due to the DRAG action on $\{\Delta,-\Delta\}$, as illustrated in FIG. 11(b). To solve this problem, we introduce the third DRAG correction with the DRAG set of $\{\alpha,\ \Delta,-\Delta\}$, then the leakage associated with the first-order process is significantly suppressed. While there is still some residual leakage error related to the second-order process on $\rm Q_{0}$ (see minor peaks in FIG. 11 with green line), but it is faint enough to be negligible. These results underscore that a driving pulse with multiple DRAG corrections can effectively archieve suppressions on both leakage and spurious transitions.
Figure 11: Suppression of more than one transitions with multiple DRAG corrections. (a) Measured sum of the excited states with $\rm Q_{0}$ when running the detection circuit with three different DRAG sets. (b) Frequency spectrum features of the driving pulses corresponding to different DRAG sets.